Find the x-intercepts
y=x^2-4x-32
Make y=0 and factor
x^2-4x-32=0
(x-8)(x+4)
Set each one to zero
x-8=0
so x=8
x+4 = 0
so x= -4

Answer by sassycarrie(1)

Make y=0 and factor
x^2-4x-32=0
(x-8)(x+4)
Set each one to zero
x-8=0
so x=8
x+4 = 0
so x= -4